If the solubility product of scandium(III) fluoride, ScF3, in pure water is 1.0 x 10-18, What is the solubility of ScF3

a. 1.0 x 10-18 M

b. 1.4 x 10-5 M

c. 6.2 10-6 M

d. 2.0 x 10-4 M

e. 2.0 10-10 M
Answer #1

SCF3ř sc 3+ + 3F [sc 3+] = abs (s be the solubility) (F) = 35 (as there are 3 Fions) Ksp = [scat] (+-) 3 = (5) (35)3 = SX 275